OpenStudy (anonymous):

(x^2-x)/(x-1) * (x+1)/(x)

jimthompson5910 (jim_thompson5910):

what do you get when you factor x^2 - x

OpenStudy (anonymous):

x

OpenStudy (shamil98):

\[\frac{ x^2-x }{ x-1 } \times \frac{ x+1 }{ x } = \frac{ x(x-1) }{ x-1 } \times \frac{ x+1 }{ x }\]

OpenStudy (shamil98):

just simplify from there.

jimthompson5910 (jim_thompson5910):

x^2 and -x have an x in common so you can factor out x to get x^2 - x x(x-1)

OpenStudy (anonymous):

Is it x+1 @shamil98

OpenStudy (shamil98):

yes.

OpenStudy (anonymous):

Thanks!